OpenStudy (anonymous):

For what value of the constant c is the function f continuous on (-infinity, infinity)?

OpenStudy (agent47):

if f(c)=(f(b)-f(a))/(b-a)

OpenStudy (agent47):

i mean if f'(c)

OpenStudy (agent47):

This is either IVT or MVT, not sure which one.

OpenStudy (agent47):

where c is between a and b and f(c) is between f(a) and f(b)

OpenStudy (anonymous):

\[f(x) = cx^2 + 2x @ x<2\]

OpenStudy (anonymous):

or f(x) = x^3-cx if x greater than or equal to two

OpenStudy (agent47):

f'(x) must be continuous, so: f'(x) = 3x^2 - c, this will be continuous for any real c.

OpenStudy (anonymous):

i was asked to find the specific value of c

OpenStudy (anonymous):

and the answer is 2/3 but i dont know how it was arrived at

OpenStudy (agent47):

hmmm.. I don't remember this then. @TuringTest

OpenStudy (turingtest):

you need to make the left and right hand limits equal at points of discontinuity

OpenStudy (turingtest):

so look for points of discontinuity in the step function, and check the limit at each junction in this case just x=2 is the only place to mess with

OpenStudy (anonymous):

\[cx^2+2x=x^3-cx\] is that right?

OpenStudy (turingtest):

at point 2, these two equations need to be equal for them to be continuous

OpenStudy (turingtest):

\[\lim_{x\to2^-}f(x)=\lim_{x\to2^+}f(x)=f(2)\]for it to be continuous

OpenStudy (anonymous):